40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Aug 23 15:24:24 2010, kiwamu, Configuration, SUS, watchdogs off 
    Reply  Mon Aug 23 18:18:22 2010, kiwamu, Configuration, SUS, Re:watchdogs off 
       Reply  Mon Aug 23 19:55:31 2010, kiwamu, Update, CDS, some progress 
Message ID: 3458     Entry time: Mon Aug 23 19:55:31 2010     In reply to: 3457
Author: kiwamu 
Type: Update 
Category: CDS 
Subject: some progress 

 {Joe, Kiwamu}

Today we did the following works in order to get ready for the new CDS test.

 - solved the DAC issue.

 - checked all the channel assignments of the ADC and the DAC.

 - preparation for modification of the AA filter chassis. 

 - checked DAC cable length.

 - connected the power cables of the BO boards to Sorensens.

 

Although we performed those works, we still couldn't do the actual damping tests.

To do the damping tests, we have to modify the AA chassis to let the SCSIs go in it. Now Joe and Steve are working for this issue.

 Also we found that we should make three more 37pin Dsub - 40pin IDC cables.  

But this is not a critical issue because the cables and the connectors are already in our hands. So we can make them any time later.

 

(Notes)


 (DAC issue)

  Now all the DAC channels are working correctly.  

There had been a combination of some issues.

  When I posted the elog entry last time, the DAC was not working at all (see here).

 But in the last week Joe found that the IO process didn't correctly run. He modified the IOP file named 'c1x02.mdl' and ran it after compiling and installing it.

 This made the situation better because we then were able to see the most of the signals coming out from the DACs.

     However we never saw any signals associated with SIDE_COILs.

 We checked the DAC cards, their slots and their timing signals. But they all were fine.

At that time we were bit confused and spent a couple of days because the DAC signals appeared at a different slot some time after we rebooted the computer. Actually this issue still remains unsolved...

Finally we found that SIDE_COILs had an input matrix which didn't show up in the medm screen.

We put 1 in the matrix and we successfully got the signal coming out from the DAC.

 

(channel assignments)

We checked all the channel assignments of the DACs and the ADCs.

All the channels are assigned correctly (i.e. pin number and channel name).

 

(AA chassis)

 We have been planning to put the SCSI cables into the AA chassis to get the ADC signals.

As Joe said in the past entry (see here) , we need a modification on the AA chassis to let the SCSIs go in it.

Joe and Steve will put an extension component so that we can make the chassis wider and eventually SCSI can go in.

 

 

(DAC cable length)

 In the default plan we are going to reuse some DAC cables which are connected to the existing systems.

To reuse them we had to make sure that the length of those cables are long enough for the new CDS.

After stopping the watchdogs, we disconnected those DAC cables and confirmed they were long enough.

Now those cables are connected to the original place where they have been before.

The same test will be performed for the binary outputs.

 

(power cables to Sorensens)

 Since the binary output boards need +/- 15V power, we hooked up the power cables to Sorensens sitting on the new 1X5 rack.

After cabling them, we turned on the power and successfully saw the green LEDs shining on the back panel of the boards.

ELOG V3.1.3-